All Exams Test series for 1 year @ ₹349 only
Question

If a random variable $x$ has the probability distribution
$x$01234567
P(x)02kk3k$2k^2$2k$k^2+k$$7k^2$

then $P(3 < x \le 6)$ is equal to

The correct answer is
0.64

Determine the value of constant k

The sum of probabilities for all possible values of a random variable must equal 1.

Given the probability distribution:

  • $P(x=0) = 0$
  • $P(x=1) = 2k$
  • $P(x=2) = k$
  • $P(x=3) = 3k$
  • $P(x=4) = 2k^2$
  • $P(x=5) = 2k$
  • $P(x=6) = k^2+k$
  • $P(x=7) = 7k^2$

Summing these probabilities gives:

$ \sum_{x=0}^{7} P(x) = 0 + 2k + k + 3k + 2k^2 + 2k + k^2 + k + 7k^2 $ $ \sum P(x) = (2+1+3+2+1)k + (2+1+7)k^2 $ $ \sum P(x) = 9k + 10k^2 $

Set the sum equal to 1:

$ 10k^2 + 9k = 1 $

Rearrange into a standard quadratic equation:

$ 10k^2 + 9k - 1 = 0 $

Factor the quadratic equation:

$ (10k - 1)(k + 1) = 0 $

This gives two possible values for $k$: $k = \frac{1}{10}$ or $k = -1$. Since probabilities must be non-negative, $k$ must be positive.

Therefore, $k = \frac{1}{10} = 0.1$.

Calculate the probability $P(3 < x \le 6)$

The required probability $P(3 < x \le 6)$ includes the values $x=4$, $x=5$, and $x=6$.

Calculate the probabilities for these values:

  • $P(x=4) = 2k^2 = 2(0.1)^2 = 2(0.01) = 0.02$
  • $P(x=5) = 2k = 2(0.1) = 0.2$
  • $P(x=6) = k^2+k = (0.1)^2 + 0.1 = 0.01 + 0.1 = 0.11$

Sum these probabilities:

$ P(3 < x \le 6) = P(x=4) + P(x=5) + P(x=6) $ $ P(3 < x \le 6) = 0.02 + 0.2 + 0.11 $ $ P(3 < x \le 6) = 0.33 $
Was this answer helpful?

Similar Questions

  1. Let $X = \{x \in \mathbb{N} : 1 \leq x \leq 19\}$ and for some $a, b \in \mathbb{R}, Y = \{ax + b : x \in X\}$. If the mean and variance of the elements of $Y$ are 30 and 750, respectively, then the sum of all possible values of $b$ is
  2. A random variable $X$ takes values $0, 1, 2, 3$ with probabilities $\frac{2a+1}{30}, \frac{8a-1}{30}, \frac{4a+1}{30}, b$ respectively, where $a, b \in \mathbf{R}$. Let $\mu$ and $\sigma$ respectively be the mean and standard deviation of $X$ such that $\sigma^2 + \mu^2 = 2$. Then $\frac{a}{b}$ is equal to :
  3. If the mean and median of the data

    x0-1010-2020-3030-4040-50 
    f362xy$\Sigma f = 20$

    are equal, then $xy^2$ is equal to

  4. If the mean of the data: 7,8,9,7,8,7,$\lambda$,8 is 8, then the variance of this data is :-

  5. Four dice are thrown simultaneously and the numbers shown on these dice are recorded in $2\times2$ matrices. The probability that such formed matrices have all different entries and are non-singular, is :

  6. Consider the following frequency distribution :

    Class0 - 66 - 1212 - 1818 - 2424 - 30
    Frequencyab1295

    If mean = $\frac{309}{22}$ and median=$14$, then the value $(a - b)^2$ is equal to ________.

  7. For 10 observations $x_1, x_2, \dots, x_{10}$, if $\sum_{i=1}^{10} (x_i + 2)^2 = 180$ and $\sum_{i=1}^{10} (x_i - 1)^2 = 90$, then their standard deviation is:
  8. The mean and variance of a data of 10 observations are 10 and 2, respectively. If an observation $\alpha$ in this data is replaced by $\beta$, then the mean and variance become 10.1 and 1.99, respectively. Then $\alpha + \beta$ equals
  9. The mean and standard deviation of 100 observations are 40 and 5.1, respectively. By mistakeone observation is taken as 50 instead of 40. If the correct mean and the correct standarddeviation are $\mu$ and $\sigma$ respectively, then $10(\mu +\sigma)$ is equal to
  10. The probability distribution of a random variable X is given below :
    X$4\text{k}$$\frac{30}{7}\text{k}$$\frac{32}{7}\text{k}$$\frac{34}{7}\text{k}$$\frac{36}{7}\text{k}$$\frac{38}{7}\text{k}$$\frac{40}{7}\text{k}$$6\text{k}$
    P(X)$\frac{2}{15}$$\frac{1}{15}$$\frac{2}{15}$$\frac{1}{5}$$\frac{1}{15}$$\frac{2}{15}$$\frac{1}{5}$$\frac{1}{15}$

    If $\text{E}(\text{X}) = \frac{263}{15}$, then $\text{P}(\text{X} < 20)$ is equal to :

Important Questions from Measures of Dispersion and Probability

  1. Let $X = \{x \in \mathbb{N} : 1 \leq x \leq 19\}$ and for some $a, b \in \mathbb{R}, Y = \{ax + b : x \in X\}$. If the mean and variance of the elements of $Y$ are 30 and 750, respectively, then the sum of all possible values of $b$ is
  2. A random variable $X$ takes values $0, 1, 2, 3$ with probabilities $\frac{2a+1}{30}, \frac{8a-1}{30}, \frac{4a+1}{30}, b$ respectively, where $a, b \in \mathbf{R}$. Let $\mu$ and $\sigma$ respectively be the mean and standard deviation of $X$ such that $\sigma^2 + \mu^2 = 2$. Then $\frac{a}{b}$ is equal to :
  3. If the mean and median of the data

    x0-1010-2020-3030-4040-50 
    f362xy$\Sigma f = 20$

    are equal, then $xy^2$ is equal to

  4. If the mean of the data: 7,8,9,7,8,7,$\lambda$,8 is 8, then the variance of this data is :-

  5. Four dice are thrown simultaneously and the numbers shown on these dice are recorded in $2\times2$ matrices. The probability that such formed matrices have all different entries and are non-singular, is :

Need Expert Advice?
More Questions from JEE Main

Start Your Preparation with Prepp Mobile App

Download the app from Google Play & App Store
Download the app from Google Play & App Store
Prepp Mobile App